OpenStudy (anonymous):

xy^2 = 4; dx/dt = -5, x=4 y=1 Evaluate dy/dt for the function at a point

OpenStudy (anonymous):

5/8

OpenStudy (anonymous):

y^2dx/dt + 2xydy/dt =0 putting the values we have -5 +8 dy/dt =0 hence dy/dt =5/8